OpenStudy (anonymous):

In Out 5 2 6 4 7 6 8 8 9 10 10 12 11 14 The table shows x-values going in and y-values coming out. The function being used is A) f(x) = 2x B) f(x) = x - 3 C) f(x) = x - 8 D) f(x) = 2x - 8

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is not \(f(x)=2x\) since that would make \(f(5)=10\) but \(f(5)=2\)

OpenStudy (misty1212):

it is not \(f(x) = x - 3 \) since that would make \(f(6)=6-3=3\) but \(f(6)=4\)

OpenStudy (anonymous):

Alright So now A. isnt right

OpenStudy (misty1212):

and it sure as hell is not \(f(x)=x-8\) since you are not subtracting 8 from the first number to get the second

OpenStudy (anonymous):

nor B right?

OpenStudy (anonymous):

Ok so then then D?

OpenStudy (misty1212):

nope not A not B not C

OpenStudy (misty1212):

gotta be D you can check to see that if you double the first number, then subtract 8 from the result, you get the second number

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Could you help with just 2 more?

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

sure why not?

OpenStudy (misty1212):

i charge $10 an answer, $20 for correct ones

OpenStudy (anonymous):

Sounds good

OpenStudy (misty1212):

just kidding...

OpenStudy (anonymous):

Solve the system. 5/8x -2/3y = -3 3/4x + 1/2y = -4 A. (9, 10) B. (9/2, 5) C. (16, -39/2) D. (-56/9, 4/3)

OpenStudy (misty1212):

lol this is donkey work, let the computer do it

OpenStudy (anonymous):

I know its not A and ok take your time

OpenStudy (misty1212):

here i will show you how

OpenStudy (misty1212):

hmm is there a typo there? none of those answers are correct

OpenStudy (anonymous):

Ill show you a picture

OpenStudy (misty1212):

\[\frac{5}{8}x-\frac{2}{3}y=-3\\ \frac{3}{4}x+\frac{1}{2}y=-4\]

OpenStudy (misty1212):

that is what i read but none of the answer choices are right yeah post a picture

OpenStudy (misty1212):

hmm that is what you wrote too, but none of those choices are correct !

OpenStudy (anonymous):

Hm idk why? well let me move on to another question (:

OpenStudy (misty1212):

hold on a sec, let me try something

OpenStudy (misty1212):

here is the right answer http://www.wolframalpha.com/input/?i=5x%2F8-2y%2F3+%3D+-3%2C3x%2F4+%2B+y%2F2+%3D+-4+

OpenStudy (anonymous):

Ok its fine!

OpenStudy (misty1212):

most of the answers to the last one don't even make logical sense, but the answer you are looking for is D an exponential function with a base bigger than one will eventually exceed any polynomial function

OpenStudy (anonymous):

Thanks so much!

OpenStudy (misty1212):

sorry about the middle one, i think they made a mistake

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Its fine (: